OpenStudy (anonymous):

What is the equation of the oblique asymptote? h(x) = x^2-3x-4/x+1 a y=x+4 b y=x^2-3 c y=x d y=x-4

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

divide

OpenStudy (misty1212):

you will get a quotient and a remainder ignore the remainder, the quotient is your obliques asymptote

OpenStudy (anonymous):

So the answer is D?

OpenStudy (anonymous):

@misty1212

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

awesome Thanks!

OpenStudy (campbell_st):

well it's not awesome as the equation doesn't have an oblique asymptote... if you factor the numerator \[\frac{x^2 - 3x -4}{x +1} = \frac{(x - 4)(x + 1)}{x +1} \] you'll find the equation simplifies to a linear form y = x - 4 and there is a point of discontinuity at x = -1 hope it helps